# Aggregation Model Hyperparameters Matter in Digital Pathology

### v2

@@ Abstract
Digital pathology has significantly advanced disease detection and pathologist efficiency through the analysis of gigapixel whole-slide images (WSI). In this process, WSIs are first divided into patches, for which a feature extractor model is applied to obtain feature vectors, which are subsequently processed by an aggregation model to predict the respective WSI label. With the rapid evolution of representation learning, numerous new feature extractor models, often termed foundational models, have emerged. Traditional evaluation methods, however, rely on fixed aggregation model hyperparameters, a framework we identify as potentially biasing the results. Our study uncovers a co-dependence between feature extractor models and aggregation model hyperparameters, indicating that performance comparability can be skewed based on the chosen hyperparameters. By accounting for this co-dependency, we find that the performance of many current feature extractor models is notably similar. We support this insight by evaluating seven feature extractor models across three different datasets with 162 different aggregation model configurations. This comprehensive approach provides a more nuanced understanding of the relationship between feature extractors and aggregation models, leading to a fairer and more accurate assessment of feature extractor models in digital pathology.
@@ 1 Introduction
The field of digital pathology (DP) has rapidly evolved, offering automated solutions for critical tasks in pathology practise, such as breast cancer [8] or metastases detection [19]. Central to this innovation is the analysis of gigapixel whole-slide images (WSI) stained with H&E. The large size of these images poses a significant computational challenge, making the direct application of standard deep learning models for classification impractical. As a solution, the digital pathology community has widely adopted an adaptation of the multiple instance learning (MIL) framework.
With this MIL approach, as depicted in step 1 in Figure 2, the WSI is partitioned into smaller, more manageable patches, known as tiles. Features from these tiles are extracted using a feature extractor model, generating an embedding vector for each tile. Within the MIL framework, all embeddings corresponding to a single WSI, collectively referred to as bag, are assigned a common label. As shown in step 2 in Figure 2, an aggregation model or MIL model is applied to the bag to learn and predict its label  [21, 11].
Figure 1: (left) Typical frameworks for evaluation of feature extraction models use fixed hyperparameters in the aggregation models, leading to substantially different results and hence limited informative value. (right) Our framework, which uses an extensive search over 81 aggregation model hyperparameters, shows a large range of results across these configurations, hence providing deeper insights into the actual performance of feature extraction models in digital pathology.
The employment of a feature extractor transcends mere computational efficiency. It is pivotal in addressing the limitations posed by scarce labeled data, a common challenge in DP. An optimal feature extractor, typically a pretrained model, generates informative features across diverse datasets. These features can be effectively utilized by a feature aggregation model, enabling efficient learning even with limited labeled data. This growing recognition of the importance of feature quality has spurred rapid progress into advanced representation learning techniques, aiming to develop more robust and informative features for digital pathology applications.
@@
Since the pivotal work of Chen et al. [6], which significantly improved visual representations using contrastive learning (SimCLR), a range of novel representation learning approaches has been introduced. SimCLR learns representations by ensuring that the embeddings of images with the same label (positive examples) are close, whereas the embeddings of images with different labels (negative examples) are far apart. Subsquently, Grill et al. [14] showed that self-supervised learning can also be done without negative examples (BYOL) and this approach was further improved and combined with transformers leading to DINO [4]. The most recent approaches combine masked autoencoder (MAE) [15] with self-distillation. This is the strategy used by iBOT [28] and is also at the core of DINOv2 [22]. The improved representations lead to impressive unsupervised learning performance on a range of natural images tasks, closing the gap to supervised learning.
@@
The adaptation of the advances in representation learning to learn better feature extractor for digital pathology has been significant. Two approaches which use contrastive learning for training the feature extractor for digital pathology are CTransPath [26] and REMEDIS [1]. Due to the large number of tiles that can be extracted from a single WSI and the availability of large publicly available datasets, such as TCGA [24], the datasets for CTransPath and REMEDIS contain 16 Mio and 50 Mio tiles, respectively. As representation learning tends to improve with a larger dataset and models, recent research efforts focus on increasing the dataset and model size while using more recent self-supervised learning methodologies. The aim of this effort is to obtain so called foundational models that are both generalizable and high performing. This would mitigate the requirement of re-training for every dataset. Filiot et al. [10] made a first step in this direction and demonstrated better classification performance compared to CTransPath by extracting feature embeddings using a model trained with iBOT. Chen et al. [5] went a step further and increased the dataset size to 100 Mio tiles while using DINOv2 to train the feature extraction model. Finally, the most recent feature extractor, Virchow [25], was also trained using the DINOv2 approach but on a dataset size of 380 Mio tiles.
Due to the rapidly increasing number of publications for feature extractor models and their rapid adoption in the field, it is now crucial to correctly compare different models. As described earlier, the feature extraction is only the first step in the classification pipeline for WSIs. Thus, when comparing feature extraction models, the second step, where the embedding vectors are processed with an aggregation model to make the final prediction, should also be accounted for. Popular model choices for this second step are AttentionMIL [16] and TransMIL [23]. Typically, when evaluating different feature extractors, an aggregation model such as AttentionMIL, is trained with a fixed set of hyperparameters for every feature extraction model. However, as illustrated in Figure 1, the ranking of the best feature extraction model drastically depends on the set of hyperparameters are used for the aggregation model training.
@@
In this paper, we show that whereas the feature extractor does indeed have some influence on the classification performance, the hyperparameter choices for the aggregation model largely dominates the results. This insight is confirmed by observing the performance fluctuation of a single feature extractor, when the aggregation model is trained with a wide range of hyperparameter sets. Thus, when comparing feature extractors, the second step of the classification pipeline, namely the aggregation model, is an important variable to control for. Our contribution in this paper is twofold.
@@
  • We reveal the interdependence between feature extractor models and aggregation model hyperparameters, challenging traditional feature extractor evaluation methods in digital pathology.
@@
Figure 2: Illustration of the typical classification pipeline with MIL in digital pathology. As a first step, features are extracted from tiles of a WSI to create embedding vectors. In a second step, the embedding vectors are aggregated using an aggregation model which outputs a prediction for the WSI. Examples of classification tasks used in this study include HER2+ vs HER2-, ABC vs GCB, and cancer vs healthy patients. In our evaluation, we test different feature extraction models while focusing on an exhaustive grid search for the aggregation model hyperparameters to see if there is a co-dependence between the two models.
2 Related Works
2.1 Feature Extraction
For the feature extraction step, it is crucial that as much information as possible is extracted of each tile, as that information can be leveraged in the subsequent classification. To extract features from the tiles, many models have been proposed. The models differ based on which dataset they are trained on, how large the model is and what learning paradigm is used. Here, we review some of the most recent and wide spread feature extraction models used in digital pathology.
ImageNet Models: ImageNet [9] has been an important benchmark dataset and many state-of-the-art models are trained on the dataset. Furthermore, the labels available for each image allow supervised training. Due to the easily available models trained on ImageNet, many digital pathology papers use them as feature extractors [20, 16]. However, the models come with the limitation that ImageNet is a natural image dataset and thus there is a domain gap to digital pathology images.
BYOL Models: The BYOL self-supervised learning approach [14] does not require negative samples, as is the case in contrastive learning, and is therefore often referred to as self-distillation [2]. This has as an important advantage for digital pathology, since labels are only assigned per WSI and are not specific to tiles. The absence of specific labels for the tiles makes BYOL one popular representation learning approach used when training feature extractors for digital pathology [12].
iBOT Models: Often, either a contrastive, self-distillation or masked autoencoder approach is used for self-supervised learning. With iBOT [28] the authors combined the masked autoencoder and self-distillation approach to benefit from each others abilities, showing strong feature extraction performance on natural images. An iBOT framework has been recently used to train a feature extractor for digital pathology [10].
DINOv2 Models: Recently, the new self-supervised learning approach DINOv2 [22] has been introduced which improves iBOT [28] by optimizing various losses and training parameters. In addition, the authors provided models which where trained on 142 Mio natural images and set a new state-of-the art for feature extraction benchmarks. Due to the success of DINOv2, several researchers have recently used the DINOv2 framework to train models on digital pathology datasets [5, 25].
CTransPath [26]: This model is trained with contrastive learning and uses a modified MoCov3 approach [7]. Often in contrastive learning for digital pathology, every tile is considered unique and all other tiles are negative examples. The authors however propose to classify similar tiles in a batch as positive examples. The authors report better classification performance using their trained feature extractor compared to using other self-supervised learning frameworks.
REMEDIS [1]: This study takes advantage of the availability of large models trained on big natural image datasets. Furthermore, these models are specifically trained for transfer learning [18]. To adapt these feature extracting models to digital pathology, the authors use the SimCLR.
2.2 Aggregation Methods
Aggregation methods have a rich literature and are continuously improving. In the context of the paper, we review two of the most widely spread approaches, namely AttentionMIL [16] and TransMIL [23]. When splitting a WSI into tiles for prediction, some assumptions need to be made about how to aggregate the predictions back to a single value that represents the whole WSI. A naive approach would be to say if any tile shows a positive label, for example indicating cancer, then the WSI should have a positive label. However, minor errors of such tile classification can easily have negative consequences. Both AttentionMIL and TransMIL use more sophisticated aggregation methods by utilizing attention with different underlying assumptions.
AttentionMIL: This method works under the assumption that, when aggregating information from the tiles, the order of the tiles does not matter and that there is no interdependency between the tiles. Nevertheless, some features from certain tiles might have a higher importance than the features of other tiles. For example, if one tile shows features related to cancer, the information of this tile should be weighted higher than when a tile shows no particular features. To this end, the authors introduce attention, which is the weight that is assigned to the features of each embedding when aggregating the features to make the final prediction for the WSI. Due to the simplicity of the approach as well as underlying assumption, this aggregation method has seen widespread adoption.
TransMIL: This method attempts to loosen the identical and independent assumption of the tile information and allow inter tile dependency. For example, some features of a tile may only become important if observed together with features from another tile. This information would be disregarded by AttentionMIL. Allowing inter tile dependency can elegantly be achieved with the self-attention mechanism of transformers. For this work, the authors use the Nyströmformer [27] architecture and show favourable results compared to several aggregation methods.
Learning Approach Model Details
Model Size
Embed.
Size
Dataset Content
Dataset Size /
Tiles (Mio.)
In-house
training
Dino v2
Combination ViT-L/16 300 1024
TCGA + in-house
35 yes
Dino v2
Combination ViT-L/14 distilled 300 1024
LVD-142M
(natural images)
142 no
iBOT Combination ViT-B 86 768 TCGA 43 no
CTransPath Contrastive Learning Swin Transformer 280 768 TCGA + PAIP 16 no
REMEDIS Contrastive Learning
BiT-M
(ResNet-152x2)
240 4096 TCGA 50 no
BYOL
(Ex Domain)
Self-distillation
ResNet50 26 2048
TCGA
2 yes
BYOL
(In Domain)
Self-distillation
ResNet50 26 2048
DLBCL, Herohe,
Camelyon16
2 yes
Table 1: List of the feature extraction models used for evaluation with some of their key properties.
@@ 3 Experimental Setup and Methods
We herein describe the pipeline employed for classification of WSI and present our hypotheses that challenge the evaluation frameworks typical in the literature. We also describe our experimental methodology to address those hypothesis, enabling a more thorough evaluation.
@@ 3.1 Pipeline for classification of WSIs
As shown in Figure 2, the typical MIL pipeline for digital pathology WSIs requires two models to obtain a final classification for a given WSI. In the first step, a feature extraction model is used. Due to the recent progress in self-supervised learning and large publicly available digital pathology datasets, these feature extraction models generalize well and are a first step towards foundational models in digital pathology [3, 25, 5]. The implication of this development means that the same feature extraction model can be used for wide range of tasks as well as different datasets. In the second step of the MIL pipeline, a smaller aggregation model is trained that takes the embedding vectors that were extracted for each tile of the WSI in the first step, and aggregates the information to make a final prediction. In contrast to the feature extraction model, this smaller aggregation model is re-trained for each dataset.
Previous studies mostly rely on frozen hyperparamters of the aggregation model when comparing feature extraction models [5, 17, 1]. This experimental setup assumes that the feature extraction model does not influence the optimal choice of aggregation model hyperparameters. Figure 1 clearly shows this assumption to be incorrect. Thus, this widely adopted evaluation framework for feature extraction models is suboptimal. Indeed, fixed aggregation model hyperparameters can favour some feature extraction models while penalizing others. We outline an experimental setup to thoroughly answer two related questions:
Question 1: Does an optimal aggregation model hyperparameter set exist, which is optimal for several feature extraction models?
Question 2: When controlling for aggregation model hyperparameters, how does the current state-of-the-art feature extraction models compare?
@@ 3.2 Feature Extraction Models
To answer these questions, we use seven different feature extraction models in our experiments. An overview of the models and their characteristics can be found in Table 1.
Firstly, we examine the potential of large models trained on natural images for the task of digital pathology feature extraction by investigating the state-of-the-art ViT-L model provided by the DINOv2 authors [22]. The model is trained on 142 Mio natural images and has 300 Mio parameters. We also investigate the performance of recently published models which are specifically trained on digital pathology datasets: CTransPath [26], REMEDIS [1] and iBOT [10]. The student model was used for iBOT as this was the default in the official codebase at the time of submission. All of these models are trained on large digital pathology datasets, as detailed in Table 1. Lastly, we investigate three feature extracting models trained in-house.
@@
Recent feature extraction models trained with the DINOv2 self-supervised learning approach for digital pathology are not available yet [25, 5]. We include this state-of-the art approach for our comparison by training a ViT-L feature extraction model from scratch. The model is trained on TCGA and a large in-house dataset with diverse tissue types and real-world data. The total training dataset size is 35 Mio $224\times 244$ tiles. WSIs are usually acquired at different magnifications. Tiles from 20$\times$ (lower) magnification WSIs have more content than those from 40$\times$ (higher) magnification. However, 40$\times$ tiles capture more details of the depicted tissue due to their higher resolution. We train on both 20$\times$ and 40$\times$ magnification, a strategy that has been shown to be beneficial [17]. We employ the official DINOv2 repository with the the default parameters for ViT-L/16 training, with a few exceptions. We decrease the batch size to 352 due to computational constraints. Due to the decreased batch size, we increase the number of epochs to 270, warm-up epochs to 25 and adjust the learning rate to $1.375\times 10^{-3}$ according to the heuristic by Goyal et al. [13].
We also train two ResNet-50 feature extractors with the BYOL self-supervised learning approach. We train one model on 2 Mio tiles, 20$\times$ magnification, which are randomly sampled from the TCGA dataset. The aim of this model is to be a lower bound of models trained on digital pathology images. The reason being that both the model size and training dataset size are smaller than other models we compare against. We refer to this model as $\text{BYOL}_{\text{ExDomain}}$ since the training dataset does not overlap with the evaluation dataset and we use it to evaluate the performance gap to larger, foundational, feature extraction models. For the second ResNet-50 model trained with BYOL, we again use a 2 Mio tile dataset. In this model, the tiles are randomly sampled from the datasets on which the models are evaluated on. Thus even though the training dataset is small compared to the other published models, there is no domain gap between the dataset on which it is trained and evaluated on. For this reason, we label the model as $\text{BYOL}_{\text{InDomain}}$. Due to the lack of a domain gap the model should also represent a higher bound for performance.
3.3 Aggregation Model Hyperparameters
To investigate the performance fluctuation of the MIL pipeline when the feature extraction model is fixed and the aggregation model hyperparameters change, we use two aggregation models, namely AttentionMIL [16] and TransMIL [23].
For each architecture, we evaluate four different hyperparameters with three values each as shown in Table 2. These are decided heuristically with preliminary experiments assessing the influence and effective range of each hyperparameter. The resulting 81 different hyperparameter configurations are outlined below for each aggregation model.
@@ 3.3 Aggregation Model Hyperparameters
AttentionMIL: When creating a batch for the aggregation model during training, there are two relevant parameters. One is the bag size, which determines the amount of tiles that is sampled from a particular WSI. The second is the bags per batch, determining how many bags from different WSIs are collected to form a batch. Here, we vary only the bag size parameter since it showed a larger influence. The Layers (fully connected layers) parameter corresponds to the architecture choice of the aggregation model. The list of numbers indicate the number of nodes for each layer. Lastly, the dropout parameter refers to the dropout which is applied at every layer of the aggregation model.
TransMIL: The selected hyperparameters for TransMIL are different due to the model architecture being a transformer and not consisting of fully connected layers. Instead of changing the number of FC layers and their respective number of nodes, for TransMIL we vary the number of Nyströmformer attention blocks. Furthermore, we also reduce the maximal bag size to 2048 due to computational limitations.
The rest of the parameters for both AttentionMIL and TransMIL training are fixed. We use a weight decay of $10^{-5}$, bags per batch of 4, AdamW optimizer, weighted cross entropy loss, and cosine annealing scheduler. We train all of the aggregation models for 50 epochs, which always converge for our hyperparameter ranges.

### v3

# The Importance of Downstream Networks in Digital Pathology Foundation Models

@@ Abstract
Digital pathology has significantly advanced disease detection and pathologist efficiency through the analysis of gigapixel whole-slide images (WSI). In this process, WSIs are first divided into patches, for which a feature extractor model is applied to obtain feature vectors, which are subsequently processed by an aggregation model to predict the respective WSI label. With the rapid evolution of representation learning, numerous new feature extractor models, often termed foundational models, have emerged. Traditional evaluation methods rely on a static downstream aggregation model setup, encompassing a fixed architecture and hyperparameters, a practice we identify as potentially biasing the results. Our study uncovers a sensitivity of feature extractor models towards aggregation model configurations, indicating that performance comparability can be skewed based on the chosen configurations. By accounting for this sensitivity, we find that the performance of many current feature extractor models is notably similar. We support this insight by evaluating seven feature extractor models across three different datasets with 162 different aggregation model configurations. This comprehensive approach provides a more nuanced understanding of the feature extractors’ sensitivity to various aggregation model configurations, leading to a fairer and more accurate assessment of new foundation models in digital pathology.
@@ 1 Introduction
Digital pathology (DP) has significantly advanced with automated solutions for tasks like breast cancer [7] and metastases detection [15], leveraging gigapixel whole-slide images (WSI) stained with H&E. The challenge of applying standard deep learning models for processing these large images has led to the adoption of the multiple instance learning (MIL) framework. In MIL, as depicted in step 1 in Figure 2, WSIs are divided into patches, also known as tiles. A feature extractor model extracts features from each tile to generate embedding vectors. These vectors, collectively referred to as a bag, are then processed by an aggregation model to predict the WSI label [16, 9]. Popular choices for aggregation models include AttentionMIL [13] and TransMIL [18], which both rely on using attention mechanisms for feature aggregation.
@@
Beyond computational efficiency, feature extractors play a critical role in overcoming the scarcity of labeled data in DP. Using representation learning approaches feature extractors can be trained on large datasets of unlabeled images enabling their use across diverse datasets. Since the pivotal work of Chen et al. [5], which significantly improved visual representations using contrastive learning (SimCLR), a range of novel representation learning approaches has been introduced. SimCLR learns representations by ensuring that the embeddings of images with the same label (positive examples) are close, whereas the embeddings of images with different labels (negative examples) are far apart. Subsquently, Grill et al. [11] showed that self-supervised learning can also be done without negative examples (BYOL). This approach was further improved and combined with transformers leading to DINO [3]. The most recent approaches combine masked autoencoder (MAE) [12] with self-distillation. This is the strategy used by iBOT [23] and is also at the core of DINOv2 [17].
@@
The DP field has adapted these representation learning advancements, notably in CTransPath [21] and REMEDIS [1]. Due to the large number of tiles that can be extracted from a single WSI and the availability of large publicly available datasets, such as TCGA [19], the datasets for CTransPath and REMEDIS contain 16 Mio and 50 Mio tiles, respectively. These large datasets allow the development of superior feature extractors, now commonly known as foundational models, that generalize across datasets without the need for re-training. Filiot et al. [8] made a first step in this direction and demonstrated better classification performance compared to CTransPath by extracting feature embeddings using a model trained with iBOT. Chen et al. [4] went a step further and increased the dataset size to 100 Mio tiles while using DINOv2 to train the feature extraction model. Finally, one of the most recent feature extractors, Virchow [20], was also trained using the DINOv2 approach but on a dataset size of 380 Mio tiles.
Foundational models promise to extract informative features from patches across diverse datasets. Ideally, capturing relevant features enhances downstream tasks, such as classification, while poor features hinder it. Feature extractors are often evaluated through their performance in basic classification tasks using models such as linear or K-NN classification [6]. However, in digital pathology, the use of an aggregation model to process embedding vectors and make final predictions can complicate the assessment of the feature extraction quality.
@@
As illustrated with an example in Figure 1, our analysis reveals that, whereas foundational models do indeed have some influence on the classification performance, they are highly sensitive to the aggregation model configuration. Thus, when comparing feature extractors, the sensitivity to the second step of the classification pipeline, namely the aggregation model, is an important variable to control for. Our contribution in this paper is twofold.
@@
  • We characterize the feature extractors’ sensitivity to various aggregation model configurations, challenging traditional feature extractor evaluation methods in digital pathology.
@@
Figure 1: Typical frameworks for evaluation of feature extraction models use fixed configurations in the aggregation models, leading to substantially different results and hence limited informative value.
@@ 2 Experimental Setup and Methods
This section outlines the classification pipeline for whole-slide images (WSIs) and outlines the framework we use for evaluating the sensitivity of feature extractors towards aggregation model configurations.
@@ 2.1 Pipeline for classification of WSIs
As depicted in Figure 2, the typical MIL pipeline for DP requires two models to obtain a final classification for a given WSI. First, a feature extraction model leverages recent self-supervised learning advancements and extensive datasets to generalize across tasks and datasets [2, 20, 4]. This model is applied to tiles in a WSI to produce feature embedding vectors. Next, a smaller aggregation model, specific to each dataset, processes the extracted embeddings to aggregate information and classify the WSI. In contrast to the feature extraction model, this aggregation model is re-trained for each dataset.
Figure 2: Illustration of the typical classification pipeline with MIL in digital pathology.
Feature extraction models have been compared under a single aggregation model configuration, i.e. fixed model architecture and hyperparameters [4, 14, 1]. Figure 1 illustrates the significant impact of aggregation model configuration choice on performance, rendering widely adopted evaluation frameworks suboptimal. Indeed, fixed aggregation model configurations can favour some feature extraction models while penalizing others. We outline an experimental setup to thoroughly address two critical questions:
Question 1: Can a single aggregation model configuration optimally support various feature extraction models?
Question 2: How do state-of-the-art feature extraction models perform relative to each other when controlling for different aggregation model configurations?
@@ 2.2 Feature Extraction Models
To explore our research questions, we assess seven feature extraction models, with details and characteristics provided in Supplementary Table S1.
We begin by evaluating the ViT-L model from DINOv2, trained on 142 million natural images with 300 million parameters, to assess the applicability of models trained on natural images for DP [17]. Additionally, we explore recently published models specifically designed for DP: CTransPath [21], REMEDIS [1], and iBOT [8] (teacher model), all trained on extensive DP datasets as detailed in Supplementary Table S1. Lastly, we investigate three feature extracting models trained in-house.
@@
We train from scratch a DINOv2 ViT-L model on TCGA and a vast in-house dataset with diverse tissue types and real-world data. The total training dataset size is 35 million $224\times 244$ tiles. WSIs are usually acquired at different magnifications. Tiles from 20$\times$ magnification offer broader content, while 40$\times$ magnification tiles provide finer tissue details due to their higher resolution. We train at both 20$\times$ and 40$\times$ magnifications to capture different features, following successful strategies in the literature [14]. We employ the official DINOv2 repository with the the default parameters for ViT-L/16 training, with a few exceptions. We decrease the batch size to 352 due to computational constraints. Due to the decreased batch size, we increase the number of epochs to 270, warm-up epochs to 25 and adjust the learning rate to $1.375\times 10^{-3}$ according to the heuristic by Goyal et al. [10].
Two ResNet-50 models are trained using BYOL: The first, $\text{BYOL}_{\text{TCGA}}$, utilizes 2 million tiles randomly sampled from the TCGA dataset at 20$\times$ magnification, featuring a smaller dataset and model size for comparison. The second, $\text{BYOL}_{\text{CHC}}$ (according to the first letter of each of the three evaluation datasets), is also trained on 2 million tiles but randomly sampled from the training set of the evaluation datasets. Thus even though the training dataset is small compared to the other published models, there is no domain gap between the dataset on which it is trained and evaluated on. This strategy ensures direct relevance to the evaluated datasets, potentially offsetting the smaller scale of the model with its domain specificity.
2.3 Aggregation Model Configurations
To investigate the performance fluctuation of the MIL pipeline when the feature extraction model is fixed and the aggregation model configuration change, we use different network hyperparameters and two well adopted aggregation model architectures: AttentionMIL [13], which uses an attention mechanism to aggregate tile information and assumes no interdependency between the tiles. TransMIL [18], which learns inter-tile dependencies by using the self-attention mechanism of transformers, in particular the Nyströmformer [22]. We change four hyperparameters with three distinct values each as shown in Table 1. These are decided heuristically with preliminary experiments assessing the influence and effective range of each hyperparameter. The resulting 162 different configurations (81 for each of the 2 architectures) are outlined below.
@@ 2.3 Aggregation Model Configurations
AttentionMIL: When creating a batch for the aggregation model during training, there are two relevant parameters. One is the bag size, which determines the amount of tiles that is sampled from a particular WSI. The second is the bags per batch, determining how many bags from different WSIs are collected to form a batch. Here, we vary only the bag size parameter since it showed a larger influence. The final batch size=bag size*bags per batch. The Layers parameter corresponds to the number of nodes in the fully connected (FC) layers in the aggregation model. The list of numbers in Table 1 indicate the number of nodes for each layer. Lastly, the dropout parameter refers to the dropout which is applied at every layer of the aggregation model.
TransMIL: The selected hyperparameters for TransMIL are different due to the model architecture being a transformer, which does not employ FC layers. Layers refers to the number of Nyströmformer attention blocks. We also reduce the maximal bag size to 2048 due to computational limitations.
Both models share fixed training parameters: a weight decay of $10^{-5}$, four bags per batch, AdamW optimizer, weighted cross entropy loss, and a cosine annealing scheduler. Aggregation models are trained for 50 epochs to ensure convergence within our configuration range.